# Dyson Equation for Correlated Linearizations and Test Error of Random Features Regression

### v2

@@ 4.2 Large Spectral Parameter
Fix $\tau\in\mathbb{R}_{\geq 0}$ and assume that $M\in\mathcal{M}$ such that, for all $z\in\mathbb{H}$, $M(z)$ solves the RDEL (5). Let $\kappa$ and $c$ be defined as in Lemma 4.7. Then, there exists $\kappa_{+}\geq\kappa+c\kappa^{-1}$ and constant $c_{+}\in\mathbb{R}_{>0}$ such that $\|\Im[M_{1,1}(z)]\|\leq c_{+}(|z|-\kappa)^{-2}(\tau+\Im[z])$ for every $z\in\{z\in\mathbb{H}\,:\,|z|\geq\kappa_{+}\}$. In particular, if $\tau=0$, then $\|\Im[M(z)]\|$ converges uniformly to 00 as $\Im[z]\downarrow 0$ on $\{z\in\mathbb{H}\,:\,\Re[z]\geq\kappa_{+}\}$.
@@
Let $m\equiv m(z)=c(|z|-\kappa)^{-1}$ be the bound in Lemma 4.7 and choose $\kappa_{+}\geq\kappa+c\kappa^{-1}$ such that $3sm^{2}\leq 2^{-1}$ and $4sm^{2}(1+2(m+m_{\star})^{2})\leq 2^{-1}$ for all $z\in\mathbb{H}$ with $|z|\geq\kappa_{+}$. Fix $z\in\mathbb{H}$ with $|z|\geq\kappa_{+}$ and denote $M=M(z)$.
By Lemma A.7, we may write $\Im[M]=M(\Im[z]\Lambda+i\tau I_{\ell}+\mathcal{S}(\Im[M]))M^{\ast}$. By Lemma 4.7, $\|M\|\leq\|M-M_{\infty}\|+\|M_{\infty}\|\leq m+m_{\star}$ where we denote $m_{\star}=\|M^{(\tau)}_{\ast}\|=\|(\mathbb{E}Q-i\tau I_{d})^{-1}\|$. Furthermore, by flatness, $\|\mathcal{S}_{1,2}(\Im[M])\|\vee\|\mathcal{S}_{2,1}(\Im[M])\|\vee\|\mathcal{S}_{2,2}(\Im[M])\|\leq s\|\Im[M_{1,1}]\|$ and $\|\mathcal{S}_{1,1}(\Im[M])\|\leq s\|\Im[M]\|$. Let $N\in\mathbb{R}^{2\times 2}$ such that $N_{j,k}=\|\Im[M_{j,k}]\|$. Then,
$N\leq\begin{bmatrix}m&m\\ m&m+m_{\star}\end{bmatrix}\begin{bmatrix}\Im[z]+\tau+s\|\Im[M]\|&s\|\Im[M_{1,1}]\|\\ s\|\Im[M_{1,1}]\|&\tau+s\|\Im[M_{1,1}]\|\end{bmatrix}\begin{bmatrix}m&m\\ m&m+m_{\star}\end{bmatrix},$
@@
$\|\Im[M_{j,k}]\|\leq m^{2}\Im[z]+2(m+m_{\star})^{2}\tau+3s(m+m_{\star})^{2}\|\Im[M_{1,1}]\|+sm^{2}\|\Im[M]\|$
for every $(j,k)\in\{(1,2),(2,1),(2,2)\}$. In particular, since $\|\Im[M]\|\leq\sum_{j,k=1}^{2}\|\Im[M_{j,k}]\|$,
$x\leq m^{2}\Im[z]+2(m+m_{\star})^{2}\tau+4s(m+m_{\star})^{2}\|\Im[M_{1,1}]\|+3sm^{2}x$
where $x=\|\Im[M_{1,2}]\|\vee\|\Im[M_{2,1}]\|\vee\|\Im[2,2]\|$. Given our choice of $\kappa_{+}$, we can rearrange to obtain $x\leq 2m^{2}\Im[z]+4(m+m_{\star})^{2}\tau+8s(m+m_{\star})^{2}\|\Im[M_{1,1}]\|$. Using the bound for $N$ above,
$\|\Im[M_{1,1}]\|$ $\leq m^{2}\Im[z]+2m^{2}\tau+4sm^{2}\|\Im[M_{1,1}]\|+m^{2}x$
@@
$+4sm^{2}(1+2(m+m_{\star})^{2})\|\Im[M_{1,1}]\|.$
Rearranging, we obtain that $\|\Im[M_{1,1}]\|\leq 2m^{2}(1+2m^{2})\Im[z]+4m^{2}(1+2(m+m_{\star})^{2})\tau$. This proves the first part of the lemma. The second part follows from the first part and the derived bound on $x$. ∎
@@ 4.3 Stieltjes Transform Representation
Given our chosen admissible set, any solution to (2) or (5) is a matrix-valued Herglotz function. Specifically, every solution possesses a Stieltjes transform representation, as guaranteed by the Nevanlinna, or Riesz-Herglotz, representation theorem [30]. This representation will prove particularly advantageous, as the positive semidefinite measure in the Stieltjes transform representation of the solution to the DEL is compactly supported.
@@
$\int_{\mathbb{R}}\Omega(\mathrm{d}\lambda)=\begin{bmatrix}I_{n}&-\mathbb{E}B^{T}(\mathbb{E}Q)^{-1}\\ -(\mathbb{E}Q)^{-1}\mathbb{E}B&(\mathbb{E}Q)^{-1}\mathbb{E}[BB^{T}](\mathbb{E}Q)^{-1}.\end{bmatrix}.$
@@
Hence, $zM_{1,1}(z)\to-I_{n}$ as $\Im[z]\to\infty$. In addition, by (14c), (14d) and Lemma 4.5, $zM_{1,2}\to\mathbb{E}B^{T}(\mathbb{E}Q)^{-1}$ and $zM_{2,1}\to(\mathbb{E}Q)^{-1}\mathbb{E}B$ as $\Im[z]\to\infty$. Finally, by (14e) and Lemma A.2,
@@
$\lim_{\Im[z]\to\infty}\mathcal{S}_{2,2}(zM)=\mathbb{E}[(B-\mathbb{E}B)\lim_{\Im[z]\to\infty}zM_{1,1}(B-\mathbb{E}B)^{T}]=-\mathbb{E}[(B-\mathbb{E}B)(B-\mathbb{E}B)^{T}].$
Since $\mathcal{S}_{2,2}(M)$ approach 00 as $\Im[z]$ approaches infinity, $(\mathcal{T}_{2,2}(M))^{-1}\to(\mathbb{E}Q)^{-1}$ as $\Im[z]\to\infty$. Also, $\mathcal{T}_{1,2}(M)\to\mathbb{E}B^{T}$ and $\mathcal{T}_{2,1}(M)\to\mathbb{E}B$ as $\Im[z]\to\infty$. We get $z(M_{2,2}-M_{\star})\to(\mathbb{E}Q)^{-1}\mathbb{E}[BB^{T}](\mathbb{E}Q)^{-1}$ as $\Im[z]\to\infty$. Since $M_{\infty}$ is real, $M-M_{\infty}$ is a matrix-valued Herglotz function. Hence, by [30, Theorem 2.3(iii) and Theorem 5.4(iv)],
@@
$\int_{\mathbb{R}}\Omega(\mathrm{d}\lambda)=-\lim_{\eta\to\infty}i\eta M(i\eta)=\begin{bmatrix}I_{n}&-\mathbb{E}B^{T}(\mathbb{E}Q)^{-1}\\ -(\mathbb{E}Q)^{-1}\mathbb{E}B&(\mathbb{E}Q)^{-1}\mathbb{E}[BB^{T}](\mathbb{E}Q)^{-1}.\end{bmatrix}$
It only remains to show that $\Omega$ is compactly supported. By Lemma 4.8, the imaginary part $\Im[M(z)]$ converges uniformly to 00 as $\Im[z]\to 0$ on $\{z\in\mathbb{H}\,:\,\Re[z]\geq\kappa_{+}\}$. Hence, by [30, Theorem 5.4(v) and Theorem 5.4(vi)], $\Omega$ is compactly supported on $[-\kappa_{+},\kappa_{+}]$, where $\kappa_{+}$ is the constant defined in Lemma 4.8. ∎
@@
It is tempting to try to directly generalize Lemma 4.9 to the solution of the regularized matrix equation (5). However, we encounter a challenge in applying the same procedure to obtain a bound on the solution to the regularized version. The issue arises from the fact that, when the regularization parameter $\tau$ is strictly positive, $M^{(\tau)}_{\infty}$ has a positive semidefinite imaginary part, which implies that the function $z\mapsto M^{(\tau)}(z)-M_{\infty}$ may not be Herglotz. One potential alternative approach is to utilize a multivariate Herglotz representation, as discussed in [49]. This representation provides an integral representation for the function $(z,i\tau)\mapsto M^{(\tau)}(z)$ involving a multivariate measure. However, it should be noted that in such representations, the measure cannot be finite unless it is trivial. Nonetheless, an analogue of Lemma 4.9 holds for the upper-left $n\times n$ block of the solution to the RDEL. The result is obtained via a similar argument, so we omit the proof.
@@
Fix $\tau\in\mathbb{R}_{\geq 0}$ and assume that $M\in\mathcal{M}$ such that, for all $z\in\mathbb{H}$, $M(z)$ solves the RDEL (5). Then, $M_{1,1}(z)=\int_{\mathbb{R}}\frac{\Omega_{1,1}(\mathrm{d}\lambda)}{\lambda-z}$ for all $z\in\mathbb{H}$, where $\Omega_{1,1}$ is a $n\times n$ matrix-valued measure on Borel subsets of $\mathbb{R}$ satisfying $\int_{\mathbb{R}}\Omega_{1,1}(\mathrm{d}\lambda)=I_{n}$.
Lemmas 4.9 and 4.11 are particularly significant because they allow us to treat the solution of the DEL as the limit of solutions to the RDEL as $\tau$ approaches zero. The tightness of the family of measures induced by the Stieltjes representation of RDEL solutions plays a key role in this step.
@@
Let $\tau\in\mathbb{R}_{>0}$. By lemmas 4.7 and 4.8, there exists $\kappa,c,\kappa_{+}\in\mathbb{R}_{>0}$ such that $\|\Im[M_{1,1}(z)]\|\leq c(|z|-\kappa)^{-2}(\tau+\Im[z])$ for every $z\in\mathbb{H}$ with $|z|\geq\kappa_{+}$. Then,
$\|\Im[M_{1,1}(\lambda+i\epsilon)]\|\leq c(\sqrt{\lambda^{2}+\epsilon^{2}}-\kappa)^{-2}(\tau+\epsilon)\leq c(\lambda-\kappa)^{-2}(\tau+\epsilon)$
for every $\lambda>\kappa_{+}$ and $\epsilon\in[0,1]$. Here, $c$ is some constant independent of $\lambda$ and $\tau$. Hence, for every $\lambda_{+}>\kappa_{+}$, according to the Stieltjes inversion formula for $\Omega^{(\tau)}$ as stated in [30, Theorem 5.4(v) and Theorem 5.4(vi)],
@@ 4.4 Power Series Representation
$(\mathbb{E}L-z\Lambda)^{-1}=\begin{bmatrix}R&-R\mathbb{E}[B^{T}]Q^{-1}\\ -Q^{-1}\mathbb{E}[B]R&Q^{-1}+Q^{-1}\mathbb{E}[B]R\mathbb{E}[B^{T}]Q^{-1}\end{bmatrix}$ (17)
with $R=(\mathbb{E}[B^{T}](\mathbb{E}Q)^{-1}\mathbb{E}[B]-\mathbb{E}[A]-zI_{n})^{-1}$. Since
$\|R\|\leq\mathrm{dist}(z,\sigma\left(\mathbb{E}[A]-\mathbb{E}[B^{T}]Q^{-1}\mathbb{E}[B]\right))^{-1},$

### v4

@@ 4.2 Large Spectral Parameter
Fix $\tau\in\mathbb{R}_{\geq 0}$ and assume that $M^{(\tau)}\in\mathcal{M}$ such that, for all $z\in\mathbb{H}$, $M^{(\tau)}(z)$ solves the RDEL (5). Let $\kappa$ and $c$ be defined as in Lemma 4.5. Then, there exists $\kappa_{+}\geq\kappa+c\kappa^{-1}$ and constant $c_{+}\in\mathbb{R}_{>0}$ such that $\|\Im[M^{(\tau)}_{1,1}(z)]\|\leq c_{+}(|z|-\kappa)^{-2}(\tau+\Im[z])$ for every $z\in\{z\in\mathbb{H}\,:\,|z|\geq\kappa_{+}\}$. In particular, if $\tau=0$, then $\|\Im[M(z)]\|$ converges uniformly to 0 as $\Im[z]\downarrow 0$ on $\{z\in\mathbb{H}\,:\,\Re[z]\geq\kappa_{+}\}$.
@@
Let $m\equiv m(z)=c(|z|-\kappa)^{-1}$ be the bound in Lemma 4.5 and choose $\kappa_{+}\geq\kappa+c\kappa^{-1}$ such that $3sm^{2}\leq 2^{-1}$ and $4sm^{2}(1+2(m+m_{\star})^{2})\leq 2^{-1}$ for all $z\in\mathbb{H}$ with $|z|\geq\kappa_{+}$. Fix $z\in\mathbb{H}$ with $|z|\geq\kappa_{+}$ and denote $M^{(\tau)}=M^{(\tau)}(z)$.
By lemmas A.6 and A.11, we may write $\Im[M^{(\tau)}]=M^{(\tau)}(\Im[z]\Lambda+i\tau I_{\ell}+\mathcal{S}(\Im[M^{(\tau)}]))(M^{(\tau)})^{\ast}$. By Lemma 4.5, $\|M^{(\tau)}\|\leq\|M^{(\tau)}-M^{(\tau)}_{\infty}\|+\|M^{(\tau)}_{\infty}\|\leq m+m_{\star}$ where we denote $m_{\star}=\|M^{(\tau)}_{\ast}\|=\|(\mathbb{E}Q-i\tau I_{d})^{-1}\|$. Furthermore, by flatness, $\|\mathcal{S}_{1,2}(\Im[M^{(\tau)}])\|\vee\|\mathcal{S}_{2,1}(\Im[M^{(\tau)}])\|\vee\|\mathcal{S}_{2,2}(\Im[M^{(\tau)}])\|\leq s\|\Im[M^{(\tau)}_{1,1}$$\|$ and $\|\mathcal{S}_{1,1}(\Im[M^{(\tau)}])\|\leq s\|\Im[M^{(\tau)}]\|$. Let $N\in\mathbb{R}^{2\times 2}$ such that $N_{j,k}=\|\Im[M^{(\tau)}_{j,k}$$\|$. Then,
$N\leq\begin{bmatrix}m&m\\ m&m+m_{\star}\end{bmatrix}\begin{bmatrix}\Im[z]+\tau+s\|\Im[M^{(\tau)}]\|&s\|\Im[M^{(\tau)}_{1,1}]\|\\ s\|\Im[M^{(\tau)}_{1,1}]\|&\tau+s\|\Im[M^{(\tau)}_{1,1}]\|\end{bmatrix}\begin{bmatrix}m&m\\ m&m+m_{\star}\end{bmatrix}$,
@@
$\|\Im[M^{(\tau)}_{j,k}]\|\leq m^{2}\Im[z]+2(m+m_{\star})^{2}\tau+3s(m+m_{\star})^{2}\|\Im[M^{(\tau)}_{1,1}]\|+sm^{2}\|\Im[M^{(\tau)}]\|$
for every $(j,k)\in\{(1,2),(2,1),(2,2)\}$. In particular, since $\|\Im[M^{(\tau)}]\|\leq\sum_{j,k=1}^{2}\|\Im[M^{(\tau)}_{j,k}$$\|$,
$x\leq m^{2}\Im[z]+2(m+m_{\star})^{2}\tau+4s(m+m_{\star})^{2}\|\Im[M^{(\tau)}_{1,1}]\|+3sm^{2}x$
where $x=\|\Im[M^{(\tau)}_{1,2}]\|\vee\|\Im[M^{(\tau)}_{2,1}]\|\vee\|\Im[M^{(\tau)}_{2,2}$$\|$. Given our choice of $\kappa_{+}$, we can rearrange to obtain $x\leq 2m^{2}\Im[z]+4(m+m_{\star})^{2}\tau+8s(m+m_{\star})^{2}\|\Im[M^{(\tau)}_{1,1}$$\|$. Using the bound for $N$ above,
$\|\Im[M^{(\tau)}_{1,1}$$\|$ $\leq m^{2}\Im[z]+2m^{2}\tau+4sm^{2}\|\Im[M^{(\tau)}_{1,1}]\|+m^{2}x$
@@
$+4sm^{2}(1+2(m+m_{\star})^{2})\|\Im[M^{(\tau)}_{1,1}$$\|$.
Rearranging, we obtain that $\|\Im[M^{(\tau)}_{1,1}]\|\leq 2m^{2}(1+2m^{2})\Im[z]+4m^{2}(1+2(m+m_{\star})^{2})\tau$. This proves the first part of the lemma. The second part follows from the first part and the derived bound on $x$. ∎
@@ 4.3 Stieltjes Transform Representation
Given our chosen admissible set, any solution to (2) or (5) is a matrix-valued Herglotz function. Specifically, every solution possesses a Stieltjes transform representation, as guaranteed by the Nevanlinna, or Riesz-Herglotz, representation theorem [matrix_herglotz]. This representation will prove particularly advantageous, as the positive semidefinite measure in the Stieltjes transform representation of the solution to the DEL is compactly supported.
@@
$\int_{\mathbb{R}}\Omega(\mathrm{d}\lambda)=\begin{bmatrix}I_{n}&-\mathbb{E}B^{\top}(\mathbb{E}Q)^{-1}\\ -(\mathbb{E}Q)^{-1}\mathbb{E}B&(\mathbb{E}Q)^{-1}\mathbb{E}[BB^{\top}](\mathbb{E}Q)^{-1}.\end{bmatrix}$.
@@
Hence, $zM_{1,1}(z)\to-I_{n}$ as $\Im[z]\to\infty$. In addition, by (14c), (14d) and Lemma 4.4, $zM_{1,2}\to\mathbb{E}B^{\top}(\mathbb{E}Q)^{-1}$ and $zM_{2,1}\to(\mathbb{E}Q)^{-1}\mathbb{E}B$ as $\Im[z]\to\infty$. Finally, by (14e) and Lemma A.1,
@@
$\lim_{\Im[z]\to\infty}\mathcal{S}_{2,2}(zM)=\mathbb{E}[(B-\mathbb{E}B)\lim_{\Im[z]\to\infty}zM_{1,1}(B-\mathbb{E}B)^{\top}]=-\mathbb{E}[(B-\mathbb{E}B)(B-\mathbb{E}B)^{\top}].$
Since $\mathcal{S}_{2,2}(M)$ approach 0 as $\Im[z]$ approaches infinity, $(\mathcal{T}_{2,2}(M))^{-1}\to(\mathbb{E}Q)^{-1}$ as $\Im[z]\to\infty$. Also, $\mathcal{T}_{1,2}(M)\to\mathbb{E}B^{\top}$ and $\mathcal{T}_{2,1}(M)\to\mathbb{E}B$ as $\Im[z]\to\infty$. We get $z(M_{2,2}-M_{\star})\to(\mathbb{E}Q)^{-1}\mathbb{E}[BB^{\top}](\mathbb{E}Q)^{-1}$ as $\Im[z]\to\infty$. Since $M_{\infty}$ is real, $M-M_{\infty}$ is a matrix-valued Herglotz function. Hence, by [matrix_herglotz, Theorem 2.3(iii) and Theorem 5.4(iv)],
@@
$\int_{\mathbb{R}}\Omega(\mathrm{d}\lambda)=-\lim_{\eta\to\infty}i\eta M(i\eta)=\begin{bmatrix}I_{n}&-\mathbb{E}B^{\top}(\mathbb{E}Q)^{-1}\\ -(\mathbb{E}Q)^{-1}\mathbb{E}B&(\mathbb{E}Q)^{-1}\mathbb{E}[BB^{\top}](\mathbb{E}Q)^{-1}.\end{bmatrix}$
It only remains to show that $\Omega$ is compactly supported. By Lemma 4.6, the imaginary part $\Im[M(z)]$ converges uniformly to 0 as $\Im[z]\to 0$ on $\{z\in\mathbb{H}\,:\,\Re[z]\geq\kappa_{+}\}$. Hence, by [matrix_herglotz, Theorem 5.4(v) and Theorem 5.4(vi)], $\Omega$ is compactly supported on $\kappa_{+},\kappa_{+}]$, where $\kappa_{+}$ is the constant defined in Lemma 4.6. ∎
@@
It is tempting to try to directly generalize Lemma 4.7 to the solution of the regularized matrix equation (5). However, we encounter a challenge in applying the same procedure to obtain a bound on the solution to the regularized version. The issue arises from the fact that, when the regularization parameter $\tau$ is strictly positive, $M^{(\tau)}_{\infty}$ has a positive semidefinite imaginary part, which implies that the function $z\mapsto M^{(\tau)}(z)-M^{(\tau)}_{\infty}$ may not be Herglotz. One potential alternative approach is to utilize a multivariate Herglotz representation, as discussed in [Luger_2017]. This representation provides an integral representation for the function $(z,i\tau)\mapsto M^{(\tau)}(z)$ involving a multivariate measure. However, it should be noted that in such representations, the measure cannot be finite unless it is trivial. Nonetheless, an analogue of Lemma 4.7 holds for the upper-left $n\times n$ block of the solution to the RDEL. The result is obtained via a similar argument, so we omit the proof.
@@
Fix $\tau\in\mathbb{R}_{\geq 0}$ and assume that $M^{(\tau)}\in\mathcal{M}$ such that, for all $z\in\mathbb{H}$, $M^{(\tau)}(z)$ solves the RDEL (5). Then, $M^{(\tau)}_{1,1}(z)=\int_{\mathbb{R}}\frac{\Omega^{(\tau)}_{1,1}(\mathrm{d}\lambda)}{\lambda-z}$ for all $z\in\mathbb{H}$, where $\Omega^{(\tau)}_{1,1}$ is a $n\times n$ matrix-valued measure on Borel subsets of $\mathbb{R}$ satisfying $\int_{\mathbb{R}}\Omega^{(\tau)}_{1,1}(\mathrm{d}\lambda)=I_{n}$.
Lemmas 4.7 and 4.8 are particularly significant because they allow us to treat the solution of the DEL as the limit of solutions to the RDEL as $\tau$ approaches zero. The tightness of the family of measures induced by the Stieltjes representation of RDEL solutions plays a key role in this step.
@@
Let $\tau\in\mathbb{R}_{>0}$. By lemmas 4.5 and 4.6, there exists $\kappa,c,\kappa_{+}\in\mathbb{R}_{>0}$ such that $\|\Im[M^{(\tau)}_{1,1}(z)]\|\leq c(|z|-\kappa)^{-2}(\tau+\Im[z])$ for every $z\in\mathbb{H}$ with $|z|\geq\kappa_{+}$. Then,
$\|\Im[M^{(\tau)}_{1,1}(\lambda+i\epsilon)]\|\leq c(\sqrt{\lambda^{2}+\epsilon^{2}}-\kappa)^{-2}(\tau+\epsilon)\leq c(\lambda-\kappa)^{-2}(\tau+\epsilon)$
for every $\lambda>\kappa_{+}$ and $\epsilon\in[0,1]$. Here, $c$ is some constant independent of $\lambda$ and $\tau$. Hence, for every $\lambda_{+}>\kappa_{+}$, according to the Stieltjes inversion formula for $\Omega^{(\tau)}$ as stated in [matrix_herglotz, Theorem 5.4(v) and Theorem 5.4(vi)],
@@ 4.4 Power Series Representation
$(\mathbb{E}L-z\Lambda)^{-1}=\begin{bmatrix}R&-R\mathbb{E}[B^{\top}]\mathbb{E}[Q]^{-1}\\ -\mathbb{E}[Q]^{-1}\mathbb{E}[B]R&\mathbb{E}[Q]^{-1}+\mathbb{E}[Q]^{-1}\mathbb{E}[B]R\mathbb{E}[B^{\top}]\mathbb{E}[Q]^{-1}\end{bmatrix}$ (17)
with $R=(\mathbb{E}[B^{\top}](\mathbb{E}Q)^{-1}\mathbb{E}[B]-\mathbb{E}[A]-zI_{n})^{-1}$. Since
$\|R\|\leq\mathrm{dist}(z,\sigma\left(\mathbb{E}[A]-\mathbb{E}[B^{\top}]Q^{-1}\mathbb{E}[B]\right))^{-1},$